simplify y= 1+2x/(1-2x)


1
Expert's answer
2022-05-16T23:37:58-0400

"y=1+\frac{2^x}{1-2^x}=\frac{1-2^x}{1-2^x}+\frac{2^x}{1-2^x}=\frac{1}{1-2^x}"